# Group Projected Subspace Pursuit for Block Sparse Signal Reconstruction: Convergence Analysis and Applications

### v1

@@ Abstract
In this paper, we present a convergence analysis of the Group Projected Subspace Pursuit (GPSP) algorithm proposed by He et al. [25] (Group Projected subspace pursuit for IDENTification of variable coefficient differential equations (GP-IDENT), Journal of Computational Physics, 494, 112526) and extend its application to general tasks of block sparse signal recovery. We prove that when the sampling matrix satisfies the Block Restricted Isometry Property (BRIP) with a sufficiently small Block Restricted Isometry Constant (BRIC), GPSP exactly recovers the true block sparse signals. When the observations are noisy, this convergence property of GPSP remains valid if the magnitude of the true signal is sufficiently large. GPSP selects the features by subspace projection criterion (SPC) for candidate inclusion and response magnitude criterion (RMC) for candidate exclusion. We compare these criteria with counterparts of other state-of-the-art greedy algorithms. Our theoretical analysis and numerical ablation studies reveal that SPC is critical to the superior performances of GPSP, and that RMC can enhance the robustness of feature identification when observations contain noises. We test and compare GPSP with other methods in diverse settings, including heterogeneous random block matrices, inexact observations, face recognition, and PDE identification. We find that GPSP outperforms the other algorithms in most cases for various levels of block sparsity and block sizes, justifying its effectiveness for general applications.
keywords:
subspace pursuit, block sparsity, feature selection.
@@ 1 Introduction
Recent advances in technology facilitate collecting and storing large amounts of high-dimensional data at low cost. Typically, each dimension represents a feature such as temperature, rating, or pricing, and many applications seek models to predict the quantity of interest based on simple combinations of these features. To achieve high accuracy while avoiding issues of over-fitting as well as redundancy, it is a common practice to focus on only a few significant features. Feature selection is a challenging problem that has been extensively studied both theoretically and numerically. Classical applications, such as DNA microarray analysis [28], image processing [13], portfolio selection [1], validate the long-lasting importance of this problem, and more recent advances in omics data analysis [46], PDE identification [40, 27, 24, 26, 25, 45, 44] and sparsity-induced deep learning [37, 35, 21, 23, 36] prove that finding critical features will become more important in the era of big data. Providing a reliable and efficient algorithm for feature selection is highly desirable in many fields.
Mathematically, the task of feature selection can be formulated as an optimization problem [6, 5]
@@
where $p\geq 1$ is usually a positive integer, $\mathbf{A}=[\mathbf{F}_{1},\dots,\mathbf{F}_{G}]\in\mathbb{R}^{N\times GM}$ with $\mathbf{F}_{g}\in\mathbb{R}^{N\times M}$ for $g=1,2,\dots,G$, $\mathbf{c}=[\mathbf{c}_{1}^{\top},\dots,\mathbf{c}_{G}^{\top}]^{\top}\in\mathbb{R}^{GM}$, $\|\mathbf{c}\|_{p,0}=\sum_{g=1}^{G}\mathcal{I}(\|\mathbf{c}_{g}\|_{p}\neq 0)$, and $\mathcal{I}$ is the indicator function that outputs $1$ when $\|\mathbf{c}_{g}\|_{p}\neq 0$ and 00 otherwise. Existence and uniqueness of solution for (1.1) are established in [9], and the counterparts for (1.2) can be found in [15, 16, 8].
Problem (1.1) and (1.2) are closely related and difficult to solve. A signal with $k$ non-zero blocks can be regarded as a signal with at most $k\times M$ non-zero entries, where $M$ denotes the size of each block. Leveraging the block sparsity pattern reduces the complexity. Indeed, the complexity of an exhaustive search for a $k$-sparse signal from solving $\mathcal{P}_{\ell_{0}}$ is $\mathcal{O}(\tbinom{M}{k}k^{2}N)$, and for a block $k$-sparse signal, this is $\mathcal{O}({\binom{G}{k}}(Mk)^{2}N)$. Moreover, we note that exactly solving (1.1) and (1.2) are both NP-hard [38, 16, 29, 14]. This means that neither of these problems has polynomial-time algorithms available. One approach to address this challenge is by convex relaxation [3, 11, 7], and the computational complexity can still be high [51].
@@
In particular, a vector $\mathbf{c}$ is called $k$-sparse if $\|\mathbf{c}\|_{0}\leq k$, and block $k$-sparse if $\|\mathbf{c}\|_{p,0}\leq k$. If $k=\|\mathbf{c}^{*}\|_{0}$, where $\mathbf{c}^{*}$ is a solution to problem (1.1), then $\mathbf{c}^{*}$ is also a solution to (1.3). Moreover, if $\mathbf{c}^{*}$ is unique, i.e., the sparsest solution [8], then by solving (1.3) with the correct $k$, we recover $\mathbf{c}^{*}$. The relation between (1.2) and (1.4) is similar. Motivated by this observation, many algorithms are developed for finding $k$-sparse or block $k$-sparse signals that best fit the data, such as alternating direction methods [12] and proximal gradient methods [4]. Among them, greedy algorithms are highly efficient and easy to implement.
@@
Iterative greedy algorithms are popular techniques for addressing (1.3) and (1.4). Some well-known methods for (1.3) include orthogonal matching pursuit (OMP) [43], subspace pursuit (SP) [10] and compressive sampling matching pursuit (CoSaMP) [39]. These algorithms find the support, i.e., indices with non-zero entries, of the unknown signal progressively by iterating different forms of candidate inclusion and exclusion. With high efficiency, their effectiveness for finding the true sparse signals is also justified by different sufficient conditions involving the restricted isometry property (RIP) [2]. For a sample matrix $\mathbf{A}\in\mathbb{R}^{N\times M}$ and a positive integer $1\leq k\leq M$, RIP specifies the bounds of eigenvalues of submatrices of $\mathbf{A}$ with $k$ columns. This is quantified by a scalar restricted isometry constant (RIC), denoted by $\delta_{k}\in(0,1)$; smaller $\delta_{k}$ implies that submatrices of $\mathbf{A}$ with $k$ columns behave more similarly as an orthonormal transformation. Based on RIP, Liu and Temlyakov [33] proved that OMP recovers a $k$-sparse signal after $k$-iterations if the sampling matrix satisfies $\delta_{k+1}<1/((1+\sqrt{2})\sqrt{k})$. This result was then sharpened by Wang and Shim [48] and Wen et al. [50]. For SP, Dai and Milenkovic [10] showed the convergence under the condition $\delta_{3k}<0.165$ which was then improved to $\delta_{3k}<0.4859$ by Song et al. [42]. Convergence analysis for CoSaMP was carried out in a series of works [39, 19, 41].
@@
These aforementioned algorithms have been extended for recovering block sparse signals as well, including block orthogonal matching pursuit (BOMP) [34], block subspace pursuit (BSP) [29] and block compressive sampling matching pursuit (BCoSaMP) [52], which are the block versions of OMP, SP and CoSaMP, respectively. The concept of RIP was extended to block RIP (BRIP) [7] for developing sufficient conditions of convergence, and they are closely related to the magnitude of the counterpart of RIC, called block Restricted Isometry Constant (BRIC). Indexed by two positive integers $M$ and $k$, a BRIC $\delta_{M,k}$ corresponding to a block matrix $\mathbf{A}$ with block size $M$ indicates the difference between any $k$ sub-blocks of $\mathbf{A}$ and an orthonormal transformation. Wang et al. [47] showed that BOMP can exactly recover the block $k$-sparse signal in $k$ steps if $\delta_{M,k+1}<1/(2\sqrt{k}+1)$, which was then improved by Wen et al. to $\delta_{M,k+1}<1/\sqrt{k+1}$ in [49]. Kamali et al. [29] proved that BSP converges to the true block $k$-sparse signal if $\delta_{M,3k}<0.1672$. As for BCoSaMP, Zhang et al. [52] showed the convergence when $\delta_{M,4k}<0.5$ is satisfied for the sampling matrix. For sufficient conditions bounding RIC and BRIC from above, it is often desirable when the respective upper bounds are closer to $1$; it allows exact signal recovery for more general sample matrices. However, it is in general difficult to evaluate the effectiveness of those sufficient conditions involving constants with different indices.
@@
The Group Projected Subspace Pursuit (GPSP) proposed in [25] was inspired by noticing limitations of state-of-the-art block-sparse signal reconstruction algorithms when applied to the identification of differential equations (DEs) with varying coefficients [27, 40, 24, 44, 45]. The objective is to identify active features governing the dynamics of observed data. Each feature corresponds to a possible differential operator depending on space and time, and the total number of true features is assumed to be small. We note that GPSP, as an extension of SP, is applicable for general block sparse regression problems. Similar to the aforementioned greedy algorithms, GPSP selects candidates by considering highly correlated features. Instead of measuring the correlation between one vector and a block of vectors by accumulating inner products, GPSP distinguishes itself by considering the projection to the column space spanned by the block of vectors. Figure 1 illustrates this key difference. Fu et al. [18] proposing BOMPR also noticed the importance of projection in block feature selection. BOMPR follows the same strategy of BOMP where candidates are not filtered once being selected; whereas GPSP incorporates the shrinking stage as BSP but with different criterion. With these algorithmic discrepancies, the convergence results established for BOMPR or BSP cannot be immediately applied, and such analysis was missing from [25]. In this paper, we investigate both theoretical and numerical aspects of GPSP to validate its effectiveness as a block-sparse signal reconstruction algorithm for general tasks.
@@
Figure 1: Two ways of measuring correlations between a vector $\mathbf{y}$ and subspaces spanned by $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$. (a) Accumulated inner product. This criterion is used in BOMP [34], BCoSaMP [52], and BSP [29]. (b) Projection to subspaces. This criterion is used in BOMPR [18] and GPSP [25].
@@
Theoretically, we establish the convergence and robustness of GPSP. Firstly, to validate GPSP as an effective greedy algorithm for block-sparse signal reconstruction, we show the exact recovery property of GPSP when the measurements are accurate and the sampling matrix satisfies BRIP with sufficiently small BRIC. See Theorem 1. Secondly, to justify the robustness of GPSP, we derive an error bound on the estimated coefficients for inaccurate measurements and prove that signals with large magnitudes can be recovered even if the data is slightly perturbed. See Theorem 2. Due to the different criteria for feature selection and removal, our proofs follow similar strategies as in [10, 29] but with new techniques for estimating changes of recovered signals during the iterations. Such analogous conditions developed for other greedy algorithms are sufficient but not necessary for exact recovery, and quantifying the BRIC for general matrices is NP-hard. In practice, it is more illuminating to compare different schemes in their algorithmic frameworks and via numerical experiments.
@@
Algorithmically, we note that GPSP share a similar structure with many state-of-the-art greedy algorithms, including BOMP [34], BOMPR [18], BCoSaMP [52], and BSP [29]. They are all iterative schemes, and during each of their respective iteration, they involve strategically expanding the pool of candidate features. For some of these algorithms, the expanding stage is followed by a shrinking stage. The expanding stage aims at including the correct features while the shrinking stage refines the selection by removing less important ones. These algorithms take different criteria for feature inclusion and exclusion. We systematically investigate such discrepancies and classify them according to the ways of quantifying correlations between residuals and matrix blocks. We also provide systematic ablation studies to compare different combinations of criteria.
@@
We present various numerical experiments to justify the superior performance of GPSP over existing methods for block sparse regression under general settings. To our best knowledge, GPSP was shown to outperform BSP in [25] for identifying PDEs with varying coefficients, and many of the existing algorithms were compared only in standard settings where each block of the sampling matrix follows an identical Gaussian distribution. In this work, we compare GPSP with state-of-the-art greedy algorithms under diverse conditions. In particular, we consider reconstructions with sampling matrices consisting of heterogeneous Gaussian blocks, where different blocks may have distinct means and variances. To examine the robustness of these algorithms against data perturbations, we conduct a series of experiments with different noise levels as well as different models of noise. Moreover, we extend the comparisons to the contexts of face recognition and PDE identification. These experiments confirm that GPSP is consistent, robust, and accurate in diverse applications with complicated block features.
@@
  • 1.
    Convergence analysis of Group Projected Subspace Pursuit (GPSP)  [25] that showed superior performances in identifying PDEs with varying coefficients. We characterize a sufficient condition for exact recovery of GPSP, and we prove an error bound for reconstruction when the observational data is perturbed.
  • 2.
@@ 2.2 Review of block restricted isometry property
In this section, we review some existing concepts used in our theoretical analysis. There are many notions developed for characterizing sufficient conditions for the exact recovery of sparse signals, such as Block Coherence (BC) [14], Mutual Subspace Incoherence (MSI) [22], and Cumulative Subspace Coherence (CSC) [16]. In this work, we develop our analysis based on the Block Restricted Isometry Property (BRIP) [15] which generalizes the well-known Restricted Isometry Property (RIP) [7].
@@
We review several properties of BRIP and BRIC that are important for subsequent analysis, and their proofs can be found in [15, 14, 29]. Intuitively, submatrices of a block matrix $\mathbf{A}$ with small BRIC are close to orthonormal matrices. This is made precise in the following lemma, which directly follows from the definition.
Lemma 1 (Equations (69) and (70) in [15]).

### v2

@@ Abstract
In this paper, we present a convergence analysis of the Group Projected Subspace Pursuit (GPSP) algorithm proposed by He et al. [HKL+23] (Group Projected subspace pursuit for IDENTification of variable coefficient differential equations (GP-IDENT), Journal of Computational Physics, 494, 112526) and extend its application to general tasks of block sparse signal recovery. We prove that when the sampling matrix satisfies the Block Restricted Isometry Property (BRIP) with a sufficiently small Block Restricted Isometry Constant (BRIC), GPSP exactly recovers the true block sparse signals. When the observations are noisy, this convergence property of GPSP remains valid if the magnitude of the true signal is sufficiently large. GPSP selects the features by subspace projection criterion (SPC) for candidate inclusion and response magnitude criterion (RMC) for candidate exclusion. We compare these criteria with counterparts of other state-of-the-art greedy algorithms. Our theoretical analysis and numerical ablation studies reveal that SPC is critical to the superior performances of GPSP, and that RMC can enhance the robustness of feature identification when observations contain noises. We test and compare GPSP with other methods in diverse settings, including heterogeneous random block matrices, inexact observations, face recognition, and PDE identification. We find that GPSP outperforms the other algorithms in most cases for various levels of block sparsity and block sizes, justifying its effectiveness for general applications.
Keywords: subspace pursuit, block sparsity, feature selection
@@ 1 Introduction
Recent advances in technology facilitate collecting and storing large amounts of high-dimensional data at low cost. Typically, each dimension represents a feature such as temperature, rating, or pricing, and many applications seek models to predict the quantity of interest based on simple combinations of these features. To achieve high accuracy while avoiding issues of over-fitting as well as redundancy, it is a common practice to focus on only a few significant features. Feature selection is a challenging problem that has been extensively studied both theoretically and numerically. Classical applications, such as DNA microarray analysis [KP07], image processing [EFM10], portfolio selection [BDDM+09], validate the long-lasting importance of this problem, and more recent advances in omics data analysis [Vin21], PDE identification [Sch17, KLL21, HKL+22, HZZ23, HKL+23, TLLK23, TLKK23] and sparsity-induced deep learning [MTK+16, MAV17, GEH19, HABN+21, Men23] prove that finding critical features will become more important in the era of big data. Providing a reliable and efficient algorithm for feature selection is highly desirable in many fields.
Mathematically, the task of feature selection can be formulated as an optimization problem [CRT06, CR07]
@@
where $p\geq 1$ is usually a positive integer, $\mathbf{A}=[\mathbf{F}_{1},\dots,\mathbf{F}_{G}]\in\mathbb{R}^{N\times GM}$ with $\mathbf{F}_{g}\in\mathbb{R}^{N\times M}$ for $g=1,2,\dots,G$, $\mathbf{c}=[\mathbf{c}_{1}^{\top},\dots,\mathbf{c}_{G}^{\top}]^{\top}\in\mathbb{R}^{GM}$, $\|\mathbf{c}\|_{p,0}=\sum_{g=1}^{G}\mathcal{I}(\|\mathbf{c}_{g}\|_{p}\neq 0)$, and $\mathcal{I}$ is the indicator function that outputs $1$ when $\|\mathbf{c}_{g}\|_{p}\neq 0$ and 00 otherwise. Existence and uniqueness of solution for (1.1) are established in [DH+01], and the counterparts for (1.2) can be found in [EM09, EV12, DE03].
Problem (1.1) and (1.2) are closely related and difficult to solve. A signal with $k$ non-zero blocks can be regarded as a signal with at most $k\times M$ non-zero entries, where $M$ denotes the size of each block. Leveraging the block sparsity pattern reduces the complexity. Indeed, the complexity of an exhaustive search for a $k$-sparse signal from solving $\mathcal{P}_{\ell_{0}}$ is $\mathcal{O}(\tbinom{M}{k}k^{2}N)$, and for a block $k$-sparse signal, this is $\mathcal{O}({\binom{G}{k}}(Mk)^{2}N)$. Moreover, we note that exactly solving (1.1) and (1.2) are both NP-hard [Nat95, EV12, KSHT13, EKB10]. This means that neither of these problems has polynomial-time algorithms available. One approach to address this challenge is by convex relaxation [CD94, Don06, CT05], and the computational complexity can still be high [YZ11].
@@
In particular, a vector $\mathbf{c}$ is called $k$-sparse if $\|\mathbf{c}\|_{0}\leq k$, and block $k$-sparse if $\|\mathbf{c}\|_{p,0}\leq k$. If $k=\|\mathbf{c}^{*}\|_{0}$, where $\mathbf{c}^{*}$ is a solution to problem (1.1), then $\mathbf{c}^{*}$ is also a solution to (1.3). Moreover, if $\mathbf{c}^{*}$ is unique, i.e., the sparsest solution [DE03], then by solving (1.3) with the correct $k$, we recover $\mathbf{c}^{*}$. The relation between (1.2) and (1.4) is similar. Motivated by this observation, many algorithms are developed for finding $k$-sparse or block $k$-sparse signals that best fit the data, such as alternating direction methods [DYZ13] and proximal gradient methods [CLK+12]. Among them, greedy algorithms are highly efficient and easy to implement.
@@
Iterative greedy algorithms are popular techniques for addressing (1.3) and (1.4). Some well-known methods for (1.3) include orthogonal matching pursuit (OMP) [TG07], subspace pursuit (SP) [DM09] and compressive sampling matching pursuit (CoSaMP) [NT09]. These algorithms find the support, i.e., indices with non-zero entries, of the unknown signal progressively by iterating different forms of candidate inclusion and exclusion. With high efficiency, their effectiveness for finding the true sparse signals is also justified by different sufficient conditions involving the restricted isometry property (RIP) [Can08]. For a sample matrix $\mathbf{A}\in\mathbb{R}^{N\times M}$ and a positive integer $1\leq k\leq M$, RIP specifies the bounds of eigenvalues of submatrices of $\mathbf{A}$ with $k$ columns. This is quantified by a scalar restricted isometry constant (RIC), denoted by $\delta_{k}\in(0,1)$; smaller $\delta_{k}$ implies that submatrices of $\mathbf{A}$ with $k$ columns behave more similarly as an orthonormal transformation. Based on RIP, Liu and Temlyakov [LT11] proved that OMP recovers a $k$-sparse signal after $k$-iterations if the sampling matrix satisfies $\delta_{k+1}<1/((1+\sqrt{2})\sqrt{k})$. This result was then sharpened by Wang and Shim [WS12] and Wen et al. [WZW+16]. For SP, Dai and Milenkovic [DM09] showed the convergence under the condition $\delta_{3k}<0.165$ which was then improved to $\delta_{3k}<0.4859$ by Song et al. [SXL14]. Convergence analysis for CoSaMP was carried out in a series of works [NT09, Fou12, SXL13].
@@
These aforementioned algorithms have been extended for recovering block sparse signals as well, including block orthogonal matching pursuit (BOMP) [LW18], block subspace pursuit (BSP) [KSHT13] and block compressive sampling matching pursuit (BCoSaMP) [ZXC+19], which are the block versions of OMP, SP and CoSaMP, respectively. The concept of RIP was extended to block RIP (BRIP) [CT05] for developing sufficient conditions of convergence, and they are closely related to the magnitude of the counterpart of RIC, called block Restricted Isometry Constant (BRIC). Indexed by two positive integers $M$ and $k$, a BRIC $\delta_{M,k}$ corresponding to a block matrix $\mathbf{A}$ with block size $M$ indicates the difference between any $k$ sub-blocks of $\mathbf{A}$ and an orthonormal transformation. Wang et al. [WLZW11] showed that BOMP can exactly recover the block $k$-sparse signal in $k$ steps if $\delta_{M,k+1}<1/(2\sqrt{k}+1)$, which was then improved by Wen et al. to $\delta_{M,k+1}<1/\sqrt{k+1}$ in [WZL+19]. Kamali et al. [KSHT13] proved that BSP converges to the true block $k$-sparse signal if $\delta_{M,3k}<0.1672$. As for BCoSaMP, Zhang et al. [ZXC+19] showed the convergence when $\delta_{M,4k}<0.5$ is satisfied for the sampling matrix. For sufficient conditions bounding RIC and BRIC from above, it is often desirable when the respective upper bounds are closer to $1$; it allows exact signal recovery for more general sample matrices. However, it is in general difficult to evaluate the effectiveness of those sufficient conditions involving constants with different indices.
@@
The Group Projected Subspace Pursuit (GPSP) proposed in [HKL+23] was inspired by noticing limitations of state-of-the-art block-sparse signal reconstruction algorithms when applied to the identification of differential equations (DEs) with varying coefficients [KLL21, Sch17, HKL+22, TLKK23, TLLK23]. The objective is to identify active features governing the dynamics of observed data. Each feature corresponds to a possible differential operator depending on space and time, and the total number of true features is assumed to be small. We note that GPSP, as an extension of SP, is applicable for general block sparse regression problems. Similar to the aforementioned greedy algorithms, GPSP selects candidates by considering highly correlated features. Instead of measuring the correlation between one vector and a block of vectors by accumulating inner products, GPSP distinguishes itself by considering the projection to the column space spanned by the block of vectors. Figure 1 illustrates this key difference. Fu et al. [FLZZ14] proposing BOMPR also noticed the importance of projection in block feature selection. BOMPR follows the same strategy of BOMP where candidates are not filtered once being selected; whereas GPSP incorporates the shrinking stage as BSP but with different criterion. With these algorithmic discrepancies, the convergence results established for BOMPR or BSP cannot be immediately applied, and such analysis was missing from [HKL+23]. In this paper, we investigate both theoretical and numerical aspects of GPSP to validate its effectiveness as a block-sparse signal reconstruction algorithm for general tasks.
@@
Figure 1: Two ways of measuring correlations between a vector $\mathbf{y}$ and subspaces spanned by $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$. (a) Accumulated inner product. This criterion is used in BOMP [LW18], BCoSaMP [ZXC+19], and BSP [KSHT13]. (b) Projection to subspaces. This criterion is used in BOMPR [FLZZ14] and GPSP [HKL+23].
@@
Theoretically, we establish the convergence and robustness of GPSP. Firstly, to validate GPSP as an effective greedy algorithm for block-sparse signal reconstruction, we show the exact recovery property of GPSP when the measurements are accurate and the sampling matrix satisfies BRIP with sufficiently small BRIC. See Theorem 1. Secondly, to justify the robustness of GPSP, we derive an error bound on the estimated coefficients for inaccurate measurements and prove that signals with large magnitudes can be recovered even if the data is slightly perturbed. See Theorem 2. Due to the different criteria for feature selection and removal, our proofs follow similar strategies as in [DM09, KSHT13] but with new techniques for estimating changes of recovered signals during the iterations. Such analogous conditions developed for other greedy algorithms are sufficient but not necessary for exact recovery, and quantifying the BRIC for general matrices is NP-hard. In practice, it is more illuminating to compare different schemes in their algorithmic frameworks and via numerical experiments.
@@
Algorithmically, we note that GPSP share a similar structure with many state-of-the-art greedy algorithms, including BOMP [LW18], BOMPR [FLZZ14], BCoSaMP [ZXC+19], and BSP [KSHT13]. They are all iterative schemes, and during each of their respective iteration, they involve strategically expanding the pool of candidate features. For some of these algorithms, the expanding stage is followed by a shrinking stage. The expanding stage aims at including the correct features while the shrinking stage refines the selection by removing less important ones. These algorithms take different criteria for feature inclusion and exclusion. We systematically investigate such discrepancies and classify them according to the ways of quantifying correlations between residuals and matrix blocks. We also provide systematic ablation studies to compare different combinations of criteria.
@@
We present various numerical experiments to justify the superior performance of GPSP over existing methods for block sparse regression under general settings. To our best knowledge, GPSP was shown to outperform BSP in [HKL+23] for identifying PDEs with varying coefficients, and many of the existing algorithms were compared only in standard settings where each block of the sampling matrix follows an identical Gaussian distribution. In this work, we compare GPSP with state-of-the-art greedy algorithms under diverse conditions. In particular, we consider reconstructions with sampling matrices consisting of heterogeneous Gaussian blocks, where different blocks may have distinct means and variances. To examine the robustness of these algorithms against data perturbations, we conduct a series of experiments with different noise levels as well as different models of noise. Moreover, we extend the comparisons to the contexts of face recognition and PDE identification. These experiments confirm that GPSP is consistent, robust, and accurate in diverse applications with complicated block features.
@@
  • Convergence analysis of Group Projected Subspace Pursuit (GPSP)  [HKL+23] that showed superior performances in identifying PDEs with varying coefficients. We characterize a sufficient condition for exact recovery of GPSP, and we prove an error bound for reconstruction when the observational data is perturbed.
@@ 2.2 Review of block restricted isometry property
In this section, we review some existing concepts used in our theoretical analysis. There are many notions developed for characterizing sufficient conditions for the exact recovery of sparse signals, such as Block Coherence (BC) [EKB10], Mutual Subspace Incoherence (MSI) [GZM09], and Cumulative Subspace Coherence (CSC) [EV12]. In this work, we develop our analysis based on the Block Restricted Isometry Property (BRIP) [EM09] which generalizes the well-known Restricted Isometry Property (RIP) [CT05].
@@
We review several properties of BRIP and BRIC that are important for subsequent analysis, and their proofs can be found in [EM09, EKB10, KSHT13]. Intuitively, submatrices of a block matrix $\mathbf{A}$ with small BRIC are close to orthonormal matrices. This is made precise in the following lemma, which directly follows from the definition.
Lemma 1 (Equations (69) and (70) in [EM09]).